OpenStudy (anonymous):

The difference of two positive numbers is 4 and the sum of their squares is 136. Find the numbers. the biggest number= the smallest number=

OpenStudy (campbell_st):

let the numbers be x and y x - y = 4 and x^2 + y^2 = 136 you need to solve simultaneously using the substitution method taking the difference equation then x = y + 4 substitute it into the squares equation (y + 4)^2 + y^2 = 136 you will have to solve a quadratic... remember the numbers are positive so only use the positive solution for y